Solve the formula for the indicated letter.
S=4Pir^2, for r^2;
P=2L + 2W, for w

Answer by stanbon(75887)

Solve the formula for the indicated letter.
S=4Pir^2, for r^2
r^2 = S/(4pi)
----------------------

P=2L + 2W, for w
2w = P-2L
w = (P-2L)/2
============
Cheers,
Stan H.